Amos has a history text in one week (seven days), but he is behind in his reading. The test will include information from 120 pages in his textbook. Amos has already read 36 pages of his textbook, and plans to read an equal amount over the course of the week. How many pages will he need to read each night to prepare for the test? Let p represent the number of pages Amos reads each night.
1
Expert's answer
2020-10-26T19:10:37-0400

if he has read 36 pages then the remaining pages are:

120 - 36 = 84


hence: he is to read 84 pages in 7 nights

let p be equal to the number of pages to read per night

the number of pages to read in 7 nights is :

p x7 = 7p


but the total number of pages to be read is 84

hence:

7p = 84

dividing both sides by 7

7p /7 = 84 /7

p =12


therefore the number of pages to be read per night is 12 pages